# LoKI: Low-damage Knowledge Implanting of Large Language Models

### v1

@@ Abstract
Fine-tuning adapts pretrained models for specific tasks but poses the risk of catastrophic forgetting (CF), where critical knowledge from pre-training is overwritten. Current Parameter-Efficient Fine-Tuning (PEFT) methods for Large Language Models (LLMs), while efficient, often sacrifice general capabilities. To address the issue of CF in a general-purpose PEFT framework, we propose Low-damage Knowledge Implanting (LoKI), a PEFT technique that is based on a mechanistic understanding of how knowledge is stored in transformer architectures. In two real-world scenarios, LoKI demonstrates task-specific performance that is comparable to or even surpasses that of full fine-tuning and LoRA-based methods across various model types, while significantly better preserving general capabilities. Our work connects mechanistic insights into LLM knowledge storage with practical fine-tuning objectives, achieving state-of-the-art trade-offs between task specialization and the preservation of general capabilities. Our implementation is publicly available as ready-to-use code [^1] https://github.com/Nexround/LoKI.
1 Introduction
Can we develop a Parameter-Efficient Fine-Tuning (PEFT) [prottasha_peft_2025, xin_parameter-efficient_2025, han_parameter-efficient_2024] method that can make the Large Language Models (LLMs) [radford2018improving, brown_language_2020] work smoothly on the target task after fine-tuning, while preventing catastrophic forgetting (CF) [DBLP:conf/iclr/KothaSR24, luo_empirical_2025]?
While PEFT techniques such as LoRA [hu2022lora] have significantly reduced the cost of adapting LLMs to downstream tasks, they often fail to preserve the model’s general capabilities. A key challenge is CF—the tendency of the model to lose general knowledge acquired during pretraining [jiang_how_2020, petroni_language_2019] when fine-tuned on specific tasks.
Recent mechanistic studies reveal that factual knowledge in LLMs associations Feed-Forward Network (FFN) layers in the transformer blocks [vaswani_attention_2017, geva_transformer_2021-1, DBLP:conf/iclr/MengSABB23]. Conventional fine-tuning approaches, however, apply indiscriminate updates across attention layers or FFN blocks, oblivious to these crucial knowledge-storing weights in general tasks. Such unguarded optimization may perturb crucial memory traces [cohen_crawling_2023, petroni_language_2019], leading to irreversible knowledge loss and degraded generalization [li_revisiting_2024, DBLP:conf/acl/HuangCWYLSYS24] - the dilemma of modern LLM adaptation.
We propose Low-damage Knowledge Implantation (LoKI), a PEFT framework that achieves state-of-the-art performance in CF mitigation while maintaining competitive task adaptation. Our approach is grounded in two key insights:
FFN layers as primary knowledge hubs. Recent studies on knowledge localization and model editing have revealed that factual and semantic knowledge in transformer models is primarily stored within the FFN layers [DBLP:conf/aaai/Li0SYMY24, DBLP:conf/iclr/MengSABB23, DBLP:conf/iclr/TanZF24, DBLP:journals/corr/abs-2310-10322, DBLP:conf/iclr/FangJWMSW0C25]. Individual weight vectors and neurons are found to express distinct facts [DBLP:conf/acl/DaiDHSCW22, geva_transformer_2022-1]. Furthermore, it has been confirmed that specific expressions of knowledge in transformer-based language models can be altered by editing neuron activations or adjusting layer weights in FFNs [DBLP:conf/acl/DaiDHSCW22, meng_locating_2022, de_cao_editing_2021-1, DBLP:conf/iclr/MengSABB23].
Redundancy in pretrained LMs. Extensive work on pruning and sparsity in language models [DBLP:conf/iclr/HeZMBN22, DBLP:conf/iclr/FrankleC19, DBLP:conf/icml/FrantarA23, DBLP:conf/nips/MichelLN19, voita_analyzing_2019] observed that LLMs harbor a substantial number of "dormant" or low-impact weights, especially in the FFN layers [DBLP:conf/iclr/KobayashiKYI24, DBLP:conf/nips/Sanh0R20, DBLP:conf/acl/ZhangL00S022]. We proved that these low-impact weights in FFN layers, located in our experiments, create safe "implant slots" where new, task-specific knowledge can be implanted without a significant impact on the general performance that LMs obtained during pre-training.
@@ 1 Introduction
By combining these two principles—allocating updates to carefully selected FFN weights—LoKI redefines the paradigm for sustainable LLM customization, ensuring models evolve without losing their essence. In summary, our main contributions are:
  • We introduce Low-damage Knowledge Implantation (LoKI), a PEFT method that simultaneously (a) adapts to downstream tasks and (b) mitigates CF in pre-trained LMs.
  • We propose Knowledge Vector Attribution (KVA), a technique to quantify the contribution of individual knowledge vectors (defined in Section 2) to the model’s stored representations. Through ablation studies, we demonstrate that KVA effectively identifies safe "implant slots" for task-specific updates.
@@
2 Knowledge in Transformer-Based Language Models
In this section, we will begin by briefly introducing the FFN layer in transformers [vaswani_attention_2017] and provide some definitions to support our discussion of the method. Following that, we will review relevant research on knowledge storage in transformer-based language models, which will establish a foundation for our proposed method.
Modern language models are built using a series of transformer layers. Each standard transformer layer contains two main components: (1) multi-head self-attention (MHSA) [vaswani_attention_2017] and (2) a position-wise feed-forward network. In this study, we focus on the FFN component, as recent research has indicated that FFNs serve as knowledge repositories within language models.
Feed-Forward Network. Consider a FFN without bias, given an input vector $x\in\mathbb{R}^{d_{\mathrm{model}}}$, the FFN applies two linear projections with a nonlinearity in between:
@@ 2 Knowledge in Transformer-Based Language Models
where $W_{\mathrm{up}}\in\mathbb{R}^{d_{\mathrm{ff}}\times d_{\mathrm{model}}}$ is the "up-projection" matrix, $W_{\mathrm{down}}\in\mathbb{R}^{d_{\mathrm{model}}\times d_{\mathrm{ff}}}$ is the "down-projection" matrix, $\sigma(\cdot)$ is an element-wise nonlinearity.
Definitions. To facilitate discussion, define the intermediate activation and the final output node:
$a=\sigma(W_{\mathrm{up}}\,x)\in\mathbb{R}^{d_{\mathrm{ff}}},\quad y=W_{\mathrm{down}}\,a\in\mathbb{R}^{d_{\mathrm{model}}}.$ (2)
Denote the $j$th output coordinate by $y_{j}=v_{j}^{\top}a,\quad v_{j}=W_{\mathrm{down},j\cdot}\in\mathbb{R}^{d_{\mathrm{ff}}}.$ We call $y_{j}$ a Knowledge Output Node and its vector $v_{j}$ a Knowledge Vector. Given the one-to-one correspondence between these two entities, we will use these terms interchangeably throughout the following sections.
FFNs as key-value memories.
FFN naturally implements a key-value memory [geva_transformer_2021-1]: each column of $W_{\mathrm{up}}$ serves as a key that detects specific input patterns, and each row of $W_{\mathrm{down}}$ acts as a value corresponding to a distinct concept in the lexical space [geva_transformer_2021-1, geva_transformer_2022-1]. This mechanism has also been validated from the perspective of gradient analysis [katz_backward_2024]. Furthermore, Geva et al. [geva_transformer_2022-1] have revealed that FFN output updates can be decomposed into multiple sub-updates induced by value vectors (knowledge vectors) in $W_{\mathrm{down}}$. Based on this insight, our focus will be exclusively on updating the parameters within $W_{\mathrm{down}}$.
Hierarchical structure of layers.
While single-layer FFNs reveal this memory behavior, deeper inspection shows a hierarchy: lower FFN layers capture surface-level patterns, whereas upper layers encode higher-level semantics [geva_dissecting_2023, DBLP:conf/acl/DaiDHSCW22, DBLP:conf/iclr/TanZF24, katz_backward_2024]. The final output distribution of transformer-based language models is gradually constructed in a bottom-up fashion [DBLP:conf/aaai/SunPNJ25, tenney_bert_2019, wallat_bertnesia_2020].
However, prior studies have demonstrated that LLMs contain a substantial number of parameters whose removal has minimal impact on the model’s performance on its primary tasks [DBLP:conf/iclr/FrankleC19, DBLP:conf/iclr/LasbyGENI24, DBLP:conf/nips/MichelLN19], particularly in FFN layers [DBLP:conf/acl/ZhangL00S022, sanh_movement_2020, DBLP:conf/icml/FrantarA23]. Based on the previous research results mentioned above, we hypothesize that these low-impact parameters in memory components can serve as "safe zones" that can be repurposed to encode new task-specific knowledge without excessively degrading the model’s original capabilities. Encouragingly, our experimental results support this hypothesis.
3 Low-damage Knowledge Implanting
Figure 1: Schematic illustration of LoKI.
@@ 3 Low-damage Knowledge Implanting
We introduce Low-damage Knowledge Implanting (LoKI), a three-stage framework—analyzing, selecting, and implanting—designed to edit transformer models with minimal disruption, thereby mitigating catastrophic forgetting (see Figure 1). Below, we briefly outline each stage before providing a detailed explanation:
@@
3.1 Analysing
We start this subsection by introducing Knowledge Vector Attribution (KVA), an approach inspired by [DBLP:conf/aaai/Hao0W021, DBLP:conf/acl/DaiDHSCW22]. Next, we will illustrate how KVA is applied in LoKI.
3.1.1 Knowledge Vector Attribution
KVA is a computational approach based on Integrated Gradients (IG) [DBLP:conf/icml/SundararajanTY17] that measures the contribution of knowledge vectors to specific output logits. To start with, we present the general formulation of IG:
[…]
where $F$ is the network function, $x^{\prime}$ is a baseline (we use $x^{\prime}=\mathbf{0}$), and $\alpha$ parameterizes the interpolation path. IG attributes predictions to input features by integrating gradients along the straight-line path from baseline $x^{\prime}$ to input $x$. To trace knowledge flow through all layers of a transformer on input sequence $\mathbf{x}$, let $\mathcal{L}$ denote the target logit, $h_{l-1}$ represent the hidden state from the previous layer. At layer $l$, the FFN first computes pre-activations
@@ 3.1.1 Knowledge Vector Attribution
We define the layer-wise, path-integrated attribution of node $j$ by
[…]
which aggregates the total gradient flowing through knowledge vector $j$ as its contribution is scaled from zero up to its actual activation. We approximate the above integral via Riemann approximation with $m$ equally-spaced steps (we set the step $m=7$ for any model):
$A_{l,j}(\mathbf{x})\approx\frac{1}{m}\sum_{k=1}^{m}\frac{\partial\mathcal{L}\bigl{(}\tfrac{k}{m}\,\mathbf{z}_{l,j}\bigr{)}}{\partial\mathbf{z}_{l,j}}.$ (6)
During each sample, KVA computes and stores $A_{l,j}(\mathbf{x})$ for every layer $l$ and every knowledge output node $j$. This yields a complete attribution log that can be analyzed post hoc to assess the contribution of each knowledge vector during inference.
3.1.2 Application to MMLU
We apply KVA on the Massive Multitask Language Understanding (MMLU) benchmark [DBLP:conf/aaai/KemkerMAHK18] to measure individual knowledge vector contributions to the pretrained model’s performance on general tasks. MMLU is ideal because:
  • it has 57 diverse subjects (STEM, humanities, professional, etc.), which covers a broad spectrum of general-knowledge tasks, ensuring extracted vectors are not domain-specialized;
@@ 3.1.2 Application to MMLU
We calculate IG scores of each FFN’s knowledge output nodes using the top 50 examples from each subject in the MMLU benchmark. We denote the recorded IG scores as $\mathbf{IG}\in\mathbb{R}^{N\times L\times D}$, where $N$ represents the number of examples, $L$ denotes the number of layers in the model, and $D$ indicates the dimensionality of the knowledge output nodes. We conducted this process using an A100 GPU for Llama3.1-8B-Instruct [DBLP:journals/corr/abs-2407-21783]. The average time required for each sample is 16.14 seconds. It’s important to note that the KVA needs to be executed only once for each model, regardless of the downstream tasks. In other words, this computational overhead occurs just once per model. Implementation details and further discussions can be found in Appendix A.
To determine if the attribution scores accurately represent the importance of individual knowledge vectors to overall task performance, we performed ablation experiments and presented the results in Section 5.1.
3.2 Selecting
As shown in Figure 2, the evaluation results of KVA reveal significant heterogeneity in the distribution of knowledge output nodes across different layers, regardless of their contribution level. Notably, both high-contribution and low-contribution nodes tend to be densely concentrated in a few middle layers and the top layers. This uneven distribution phenomenon among nodes with high contributions aligns with the findings of Meng et al. [meng_locating_2022, DBLP:conf/iclr/MengSABB23] and Dai et al. [dai_knowledge_2022-1]. Our work further observed the phenomenon of uneven distribution of low contribution nodes.
Given the established relationships between the layers within transformers [sun2025transformer, geva_transformer_2022-1], focusing weight updates on a few layers for downstream task adaptation may disrupt this hierarchical relationship. Our experiments revealed that this uneven parameter allocation will lead to a more pronounced phenomenon of catastrophic forgetting (see Section 5.2 for details). Additionally, Hase et al. [DBLP:conf/nips/HaseBKG23] have found a significant challenge: even when factual knowledge resides in specific layers, such as the mid-layer FFNs, modifying weights in distant layers, especially the earlier ones, can effectively "override" the information flow downstream.
@@ 3.2 Selecting
Figure 2: Heatmaps of top $5\%$ KVA results across all 32 FFN layers of Llama3.1-8B-Instruct. The upper (red-tinted) map highlights the distribution of high-contribution node positions, while the lower (blue-tinted) map highlights the distribution of low-contribution node positions. Color intensity (log-scale) indicates the density of nodes in each category, with darker colors representing higher density. Heatmaps of other models can be found in Appendix B.
3.2.1 Layer-Balanced Strategy
Based on the earlier considerations, we propose the Layer-Balanced Strategy, which is a strategy that determines the trainable parameter positions under the constraint of allocating the same number of parameters to each layer. This strategy aims to (a) ensure that newly implanted knowledge conforms to the hierarchical relationship of the transformer model’s knowledge structure, and (b) avoid disproportionately disturbing its inherent knowledge hierarchy.
@@ 3.2.1 Layer-Balanced Strategy
Let the model have $L$ layers, where only parameters within each layer’s $W_{\mathrm{down}}$ matrix are partially trainable, while all other parameters across all layers remain frozen. Each $W_{\mathrm{down}}$ in layer $l$ contains $D_{l}$ knowledge output nodes (with $D_{l}=D$ assumed identical across layers for simplicity). The hyperparameter $q\in(0,100)$ governs the percentage of trainable nodes selected from all $W_{\mathrm{down}}$ matrices. Given $N$ inference samples, we denote the integrated gradients for node $i$ in layer $l$ on sample $t$ as:
$\mathbf{IG}_{l,i}^{(t)},\quad i=1,\dots,D,\;t=1,\dots,N$
@@
    Per-Sample Normalization & Local Selection: For each sample $t$ and layer $l$, normalize gradients:
    $\hat{\mathbf{IG}}_{l,i}^{(t)}=\frac{\mathbf{IG}_{l,i}^{(t)}-\min_{j}\mathbf{IG}_{l,j}^{(t)}}{\max_{j}\mathbf{IG}_{l,j}^{(t)}-\min_{j}\mathbf{IG}_{l,j}^{(t)}}$ (8)
    Select $k_{l}$ nodes with the smallest normalized values:
    $S_{l}^{(t)}=\underset{i\in\{1,\dots,D\}}{\mathrm{argsort}\downarrow}\left(\hat{\mathbf{IG}}_{l,i}^{(t)}\right)[1:k_{l}]$ (9)

### v2

@@ Abstract
Fine-tuning adapts pretrained models for specific tasks but poses the risk of catastrophic forgetting (CF), where critical knowledge from pretraining is overwritten. To address the issue of CF in a general-purpose framework, we propose Low-damage Knowledge Implanting (LoKI), a parameter-efficient fine-tuning (PEFT) technique that utilizes recent mechanistic understanding of how knowledge is stored in transformer architectures. We compare LoKI against state-of-the-art PEFT methods in two real-world fine-tuning scenarios. The results show that LoKI demonstrates significantly better preservation of general capabilities. At the same time, its task-specific performance is comparable to or even surpasses that of full parameter fine-tuning and these PEFT methods across various model architectures. Our work bridges the mechanistic insights of LLMs’ knowledge storage with practical fine-tuning objectives, enabling an effective balance between task-specific adaptation and the retention of general-purpose capabilities.
Codehttps://github.com/Nexround/LoKI
Extended versionhttps://arxiv.org/abs/2505.22120
Introduction
Transformer-based language models (vaswani_attention_2017; petroni_language_2019) accumulate extensive world knowledge during pretraining (Radford2018ImprovingLU; brown_language_2020), which becomes implicitly embedded in their parameters. Owing to their broad generalization capabilities, they serve as a strong foundation for downstream task adaptation, prompting significant efforts to develop increasingly efficient fine-tuning methods (prottasha_peft_2025; xin_parameter-efficient_2025). However, the process of fine-tuning pretrained models for downstream tasks is often accompanied by catastrophic forgetting (CF) (DBLP:conf/iclr/KothaSR24; luo_empirical_2025), a phenomenon where the model loses previously acquired capabilities after fine-tuning. Recent studies have pointed out that conventional fine-tuning approaches (prottasha_peft_2025; xin_parameter-efficient_2025) typically perform indiscriminate updates across modules within transformer architectures, oblivious to these crucial knowledge-storing weights in general tasks. Such unconstrained optimization may perturb crucial memory traces (cohen_crawling_2023; petroni_language_2019), leading to irreversible knowledge loss and degraded generalization performance (li_revisiting_2024; DBLP:conf/acl/HuangCWYLSYS24). Extensive research has investigated the internal structure of large language models (LLMs) (DBLP:conf/aaai/Li0SYMY24; DBLP:conf/iclr/MengSABB23), revealing that model-internal knowledge is both identifiable and editable via targeted interventions such as knowledge localization and rewriting (geva_transformer_2021-1; DBLP:conf/iclr/MengSABB23). Moreover, the inherent redundancy in LLMs has been shown to support techniques like sparsification and parameter pruning (DBLP:conf/iclr/HeZMBN22; DBLP:conf/iclr/FrankleC19), enabling more efficient representation without significantly degrading performance (DBLP:conf/iclr/LasbyGENI24). However, these findings have yet to be effectively integrated into the field of model fine-tuning.
Figure 1: Schematic illustration of the staged fine-tuning process in LoKI.
These findings lead us to hypothesize that it is possible to identify low-contributing weights in pretrained models, which can then serve as capacity for injecting new, task-specific knowledge with minimal impact on existing competencies. Based on this assumption, we propose Low-damage Knowledge Implanting, as LoKI, a parameter-efficient fine-tuning method that leverages insights from interpretability studies on the internal knowledge storage mechanisms of LLMs. Aiming to mitigate catastrophic forgetting, LoKI consists of three stages: analyzing, selecting, and implanting (see Figure 1). In the analyzing stage, we introduce Knowledge Vector Attribution (KVA), a gradient-based attribution technique (DBLP:conf/icml/SundararajanTY17) that evaluates the contribution of each vector in the down-projection matrix $W_{\mathrm{down}}$ (see Background section for details) to the model’s pretrained behavior. In the selecting stage, motivated by the known interdependence between transformer layers (DBLP:conf/aaai/SunPNJ25)—especially in the progressive refinement of knowledge (geva_dissecting_2023; DBLP:conf/acl/DaiDHSCW22)—we propose the Layer-Balanced Strategy, which ensures that new knowledge aligns with the model’s hierarchical structure. Leveraging KVA results, this strategy enforces an equal quota of trainable parameters per layer by decomposing each $W_{\mathrm{down}}$ into two subsets: $W_{\mathcal{S}}$ (active) and $W_{\setminus\mathcal{S}}$ (frozen). This yields a layer-balanced trainable subset $\mathbb{W}_{\mathcal{S}}$. In the implanting stage, we freeze all model parameters except $\mathbb{W}_{\mathcal{S}}$, which is updated via fine-tuning. Compared to existing fine-tuning approaches, LoKI offers several major advantages.
  • Superior balance between CF and task-specific performance. Fine-tuning LLMs of two representative sizes with LoKI effectively mitigates CF, outperforming state-of-the-art PEFT methods while maintaining strong task performance.
  • Intrinsically parameter-efficient. LoKI updates only a subset of the model’s original parameters and allows for explicit control over the number of trainable weights.
  • Synergistic with other tuning methods. In addition to adapting directly for downstream tasks, we demonstrate that LoKI can be combined with existing parameter-efficient tuning techniques such as Low-Rank Adaptation (LoRA) (hu2022lora), further reducing the number of trainable parameters.
@@ Introduction
By allocating updates to carefully selected weights, LoKI provides a competitive approach to sustainable LLM customization, enabling models to evolve while preserving their core capabilities. In summary, our main contributions are:
  • We introduce LoKI, a fine-tuning framework for LLMs that primarily aims to mitigate CF during adaptation to downstream tasks.
  • We propose KVA, a technique to quantify the contribution of individual knowledge vectors (defined in the Background section) to the model’s stored representations. Furthermore, using this method, we uncover a surprising phenomenon: in transformer models, knowledge vectors with both globally high and low contributions tend to be densely located in similar layers.
@@
Related Works
Catastrophic Forgetting
CF has long been recognized as a critical challenge in neural networks (DBLP:conf/nips/French93; kemker2018measuring), and recent studies have begun to analyze its manifestation specifically in LLMs (li_revisiting_2024; DBLP:conf/iclr/KothaSR24). Representative methods include: Replay-based methods (DBLP:conf/nips/dAutumeRKY19; DBLP:conf/nips/Lopez-PazR17; DBLP:conf/iclr/ChaudhryRRE19) interleave data from previous tasks during training to reinforce prior knowledge. Regularization-based methods (DBLP:journals/pami/LiH18a; DBLP:journals/corr/KirkpatrickPRVD16; deng_flattening_2021) constrain parameter updates to prevent drift from previously important weights. However, most of the previous methods generally treat knowledge retention as a black box, rarely considering how pretrained LLMs organize and store knowledge internally.
Parameter‑Efficient Fine‑Tuning
Recent advances in PEFT aim to mitigate CF by preserving pretrained knowledge while enabling task adaptation. For example, DBLP:conf/icml/Zhu [^0] YD0K24 refines a small subset of critical parameters to retain performance on original tasks. However, it requires task-specific parameter selection. Some methods structurally decouple knowledge to resist CF. CorDA (DBLP:conf/nips/YangLZSWNG24) freezes dominant singular directions, assumed to encode general knowledge, while adapting residual subspaces. LoRASculpt (DBLP:conf/cvpr/0001HWYY25) uses magnitude-based masking and conflict-aware regularization to constrain LoRA (hu2022lora) updates within critical knowledge regions. Orthogonal subspace approaches like O-LoRA (DBLP:conf/emnlp/WangCGXBZZGH23) and LoRI (DBLP:journals/corr/abs-2504-07448) reduce task interference by isolating updates in mutually orthogonal directions, effective in multi-task scenarios but less focused on preserving original capabilities. In contrast, LoKI proposes a one-time knowledge localization strategy, directly quantifying parameter contributions to general knowledge, enabling scalable and effective CF resistance across diverse downstream tasks.
Knowledge Locating and Editing
ROME (meng_locating_2022) revises individual factual associations by directly modifying the FFN weight vectors. KN (DBLP:conf/acl/DaiDHSCW22) regulates the expression of specific facts by controlling the activation levels of identified neurons. AlphaEdit (DBLP:conf/iclr/FangJWMSW0C25) further advances this line by applying zero-space projection to the output weight matrix of FFN layers to perform targeted edits. However, despite their effectiveness in editing specific knowledge expressions, these methods have yet to be seamlessly integrated into parameter-efficient fine-tuning pipelines that aim to retain the model’s broad pretrained capabilities. Moreover, compared to existing knowledge localization methods, KVA and the Layer-Balanced Strategy constitute a complete model analysis pipeline in practice. This combination transcends the limitation of focusing solely on specific factual expressions, evolving the underlying ideas into a general-purpose approach for model analysis.
Background
We will begin this section by briefly reviewing transformer architecture and relevant research on knowledge storage in transformer-based language models, which will establish a foundation for our proposed method. Transformer-based language models are built from stacked layers, each consisting of two primary components: a multi-head self-attention (MHSA) mechanism (vaswani_attention_2017) and a position-wise feed-forward network (FFN). While some studies have explored the role of MHSA in transformers (voita_analyzing_2019; DBLP:conf/blackboxnlp/ClarkKLM19; DBLP:conf/blackboxnlp/VigB19), a growing number of studies focus on the FFN layers for knowledge localization and editing (geva_transformer_2021-1; geva_transformer_2022-1; katz_backward_2024). For an input vector $x\in\mathbb{R}^{d_{\mathrm{model}}}$, a typical FFN without bias applies two linear transformations with a non-linearity in between:
@@
where $W_{\mathrm{up}}\in\mathbb{R}^{d_{\mathrm{ff}}\times d_{\mathrm{model}}}$ is the up-projection matrix, $W_{\mathrm{down}}\in\mathbb{R}^{d_{\mathrm{model}}\times d_{\mathrm{ff}}}$ is the down-projection matrix, and $\sigma(\cdot)$ denotes an element-wise activation function. Recent studies suggest that FFN layers function as linear associative memories, where the $W_{\mathrm{up}}$ acts as a collection of keys detecting input patterns, and the $W_{\mathrm{down}}$ contains values corresponding to interpretable concepts (geva_transformer_2021-1; geva_transformer_2022-1). This view has been reinforced by both activation- and gradient-based analyses (katz_backward_2024). Specifically, each output coordinate $y_{j}$ in the FFN can be viewed as a knowledge output node, computed as $y_{j}=v_{j}^{\top}\sigma(W_{\mathrm{up}},x)$, where $v_{j}$ is the $j$-th row of $W_{\mathrm{down}}$ and serves as a knowledge vector. On the other hand, while single-layer FFN reveals this memory behavior, deeper inspection shows a hierarchy: lower FFN layers capture surface-level patterns, whereas upper layers encode higher-level semantics (geva_dissecting_2023; DBLP:conf/acl/DaiDHSCW22; DBLP:conf/iclr/TanZF24; katz_backward_2024). The final output distribution of transformer-based language models is gradually constructed in a bottom-up fashion (DBLP:conf/aaai/SunPNJ25; tenney_bert_2019; wallat_bertnesia_2020). Such insights have inspired a wave of model editing techniques that target the FFN layers—particularly $W_{\mathrm{down}}$—to insert, update, or erase factual knowledge without extensive retraining (geva_transformer_2022-1; dai_knowledge_2022-1; meng_mass-editing_2023). Motivated by the success of these approaches, our study focuses on implanting task-specific knowledge through FFNs, which we consider the main site for implanting new knowledge into the model. Importantly, prior research reveals that LLMs exhibit considerable parameter redundancy, especially within FFNs (DBLP:conf/iclr/KobayashiKYI24; DBLP:conf/nips/Sanh0R20; DBLP:conf/acl/ZhangL00S022; sanh_movement_2020; DBLP:conf/icml/FrantarA23). Numerous studies show that substantial portions of model weights can be pruned or restructured with minimal performance loss on general tasks (DBLP:conf/iclr/FrankleC19; DBLP:conf/nips/MichelLN19; sanh_movement_2020; DBLP:conf/icml/FrantarA23). Based on the previous research mentioned above, we hypothesize that these low-impact parameters in FFN layers can be repurposed to encode new task-specific knowledge without excessively degrading the model’s original capabilities. Our experimental results provide evidence supporting this hypothesis.
Low-damage Knowledge Implanting
@@
We introduce LoKI, a three-stage framework—analyzing, selecting, and implanting—designed to edit transformer models with minimal disruption, thereby mitigating CF (see Figure 1). Below, we briefly outline each stage before providing a detailed explanation:
@@
Analysing
We start this subsection by introducing Knowledge Vector Attribution (KVA), an attribution method inspired by (DBLP:conf/aaai/Hao0W021; DBLP:conf/acl/DaiDHSCW22), designed to evaluate the contribution of individual knowledge vectors to the model’s performance. Next, we illustrate the workflow of the analysis stage—in other words, how KVA is utilized within the LoKI framework.
Knowledge Vector Attribution
KVA is a computational approach based on Integrated Gradients (IG) (DBLP:conf/icml/SundararajanTY17) that measures the contribution of knowledge vectors to specific output logits.
$\mathbf{IG}_{i}(x)=(x_{i}-x^{\prime}_{i})\int_{0}^{1}\frac{\partial F(x^{\prime}+\alpha(x-x^{\prime}))}{\partial x_{i}}d\alpha,$ (2)
where $F$ is the network function, $x^{\prime}$ is a baseline (we use $x^{\prime}=\mathbf{0}$), and $\alpha\in[0,1]$ is an interpolation coefficient that defines the integration path from baseline to input. IG attributes predictions to input features by integrating gradients along the straight-line path from baseline $x^{\prime}$ to input $x$. To trace knowledge flow through all layers of a transformer on input sequence $\mathbf{x}$, let $\mathcal{L}$ denote the target logit, $h_{l-1}$ represent the hidden state from the previous layer. At layer $l$, the FFN first computes pre-activations
@@
As mentioned above, we treat FFNs as collections of key-value memory pairs. To attribute the contribution of each knowledge vector to the final output of an LLM, we define the layer-wise, path-integrated attribution of node $j$ by
$Attr_{l,j}(\mathbf{x})=\int_{0}^{1}\frac{\partial\mathcal{L}\bigl(\alpha\,\mathbf{z}_{l,j}\bigr)}{\partial\mathbf{z}_{l,j}}\,d\alpha,$ (4)
which aggregates the total gradient flowing through knowledge vector $j$ as its contribution is scaled from zero up to its actual activation. We approximate the above integral via Riemann approximation with $m$ equally-spaced steps (we set the step $m=7$ for any model, and further discussion on the setting of $m$ can be found in Appendix A):
$Attr_{l,j}(\mathbf{x})\approx\frac{1}{m}\sum_{k=1}^{m}\frac{\partial\mathcal{L}\bigl(\tfrac{k}{m}\,\mathbf{z}_{l,j}\bigr)}{\partial\mathbf{z}_{l,j}}.$ (5)
During each sample, KVA computes and stores $Attr_{l,j}(\mathbf{x})$ for every layer $l$ and every knowledge output node $j$. This yields a complete attribution log that can be analyzed post hoc to assess the contribution of each knowledge vector during inference.
Scoring Knowledge Vectors
Pretrained LLMs exhibit a wide range of capabilities, including world knowledge, common-sense reasoning, and instruction following. While it is difficult to precisely determine the full scope of an LLM’s internal knowledge, it can be approximated through extensive textual input (DBLP:conf/iclr/MengSABB23). To evaluate the contribution of individual knowledge vectors to general-purpose performance—and to identify those with minimal impact on core capabilities—a critical factor is the choice of dataset for quantitative analysis. To ensure broad domain coverage, we apply KVA to the Massive Multitask Language Understanding (MMLU) benchmark (DBLP:conf/aaai/KemkerMAHK18). MMLU is ideal because:
  • it has 57 diverse subjects (STEM, humanities, professional, etc.), which covers a broad spectrum of general-knowledge tasks, ensuring the evaluation score of vectors is not domain-specialized;
@@
This process, executed on an RTX4090 GPU using Llama3.1-8B-Instruct (DBLP:journals/corr/abs-2407-21783), takes an average of 9.69 seconds per sample. In parallel, we apply KVA to the full MMLU dataset. Results show that the final selected nodes—based on the Layer-Balanced Strategy (detailed later)—overlap by 97.57% between the reduced set and the full dataset. This confirms that our sampling strategy substantially reduces computational cost while preserving high fidelity in attribution outcomes. Notably, KVA needs to be performed only once per model, independent of downstream tasks. That is, this computational overhead is incurred a single time per model. Implementation details and additional discussion are provided in Appendix A.
Selecting
@@
Figure 2: Heatmaps of the top $5\%$ KVA results across all 32 layers of Llama3.1-8B-Instruct. The vertical axis denotes node indices, and the horizontal axis denotes layer indices. The upper (red-tinted) heatmap illustrates the distribution of high-contribution node positions, while the lower (blue-tinted) heatmap illustrates the distribution of low-contribution node positions. Color intensity (log-scale) reflects the density of nodes within each category, with darker colors indicating higher density. Heatmaps for additional models are provided in Appendix B.
As shown in Figure 2, the evaluation results of KVA reveal significant heterogeneity in the distribution of knowledge output nodes across different layers, regardless of their contribution level. This uneven distribution phenomenon among nodes with high contributions aligns with the findings of Meng et al. (meng_locating_2022; DBLP:conf/iclr/MengSABB23) and Dai et al. (dai_knowledge_2022-1). Notably, our results uncover a novel phenomenon: both high- and low-contribution nodes are densely concentrated in the same layers. This suggests a potential structural relationship in the placement of knowledge vectors across different attribution levels. We believe this observation may offer new insights into inter-layer knowledge organization in transformer architectures. As our current focus is on leveraging the analysis results for guiding fine-tuning, we leave a deeper investigation of this phenomenon to future work. These findings shed a critical light on fine-tuning strategies: the layer-wise distribution of knowledge is not random but structurally biased. Consequently, naïve parameter-efficient approaches that update only a limited number of layers may inadvertently disrupt the intrinsic knowledge hierarchy of transformers (sun_transformer_2025; geva_transformer_2022-1). Our experiments show that such imbalanced parameter allocation exacerbates catastrophic forgetting (see the Ablation Studies section). Furthermore, Hase et al. (DBLP:conf/nips/HaseBKG23) have found a significant challenge: even when factual knowledge resides in specific layers, such as the mid-layer FFNs, modifying weights in distant layers, especially the earlier ones, can effectively ”override” the information flow downstream.
Layer-Balanced Strategy
@@
Based on the previous considerations, we propose the Layer-Balanced Strategy, which is a strategy that determines the trainable parameter positions under the constraint of allocating the same number of parameters to each layer. This strategy aims to (a) ensure that newly implanted knowledge conforms to the hierarchical relationship of the transformer model’s knowledge structure, and (b) avoid disproportionately disturbing its inherent knowledge hierarchy. Let the model have $L$ layers, where only parameters within each layer’s $W_{\mathrm{down}}$ matrix are partially trainable, while all other parameters across all layers remain frozen. Each $W_{\mathrm{down}}$ in layer $l$ contains $D_{l}$ knowledge output nodes (with $D_{l}=D$ assumed identical across layers for simplicity). The hyperparameter $q\in(0,100)$ governs the percentage of trainable nodes selected from all $W_{\mathrm{down}}$ matrices. Given $N$ inference samples, we denote the KVA result for node $i$ in layer $l$ on sample $t$ as:
$\mathbf{Attr}_{l,i}^{(t)},\quad i=1,\dots,D,\;t=1,\dots,N$
@@
    Per-Sample Local Selection: For each sample $t$ and layer $l$, select $k_{l}$ nodes with the smallest values:
    $S_{l}^{(t)}=\underset{i\in\{1,\dots,D\}}{\mathrm{argsort}\downarrow}\left(\hat{\mathbf{Attr}}_{l,i}^{(t)}\right)[1:k_{l}]$ (7)